OpenStudy (anonymous):

Jodie bicycles 8 km/c faster than Javier. In the same time it takes Javier to bicycle 57 km, Jodie can bicycle 81 cm. How fast does each bicylist travel?

OpenStudy (anonymous):

suppose the speed of jevier's cycle is x. hence speed of jodie's cycle = x+8. => 57/x=81/(x+8) => 81x=57x + 8*57 => 81x-57x=8*57 => 24x=8*57 => x=19 x+8= 19+8=27.

OpenStudy (amistre64):

81 centimeters?

OpenStudy (turingtest):

correct, I hate trying to use the math app sometimes...

OpenStudy (turingtest):

I assumed she meant Km

OpenStudy (anonymous):

I need cm for both Javier and Jodie.

OpenStudy (anonymous):

OH...crud...KM for both.